OpenStudy (anonymous):

(-4,2) and (-1,-2) find the slope

OpenStudy (anonymous):

4/3

OpenStudy (anonymous):

Would I Be -4/3 ?!

OpenStudy (anonymous):

yean negative

OpenStudy (anonymous):

can someone help me please

OpenStudy (anonymous):

(-1,2) and (2,-3) find the slope.... I got -5/3 ?! Is it right?? @Michele_Laino

OpenStudy (anonymous):

let me check @Juliannaaaaaa

OpenStudy (anonymous):

K :)

OpenStudy (michele_laino):

here we have to apply the same formula: \[\Large m = \frac{{{y_2} - {y_1}}}{{{x_2} - {x_1}}} = \frac{{ - 3 - 2}}{{2 - \left( { - 1} \right)}} = ...?\]

OpenStudy (anonymous):